### v1

# Compact Speech Translation Models via Discrete Speech Units Pretraining

Abstract
Using Self-Supervised Learning (SSL) as model initialization is now common to obtain strong results in Speech Translation (ST). However, they also impose a large memory footprint, hindering on-device deployment. In this paper, we leverage the SSL models by pretraining smaller models on their Discrete Speech Units (DSU). We pretrain encoder-decoder models on 1) Filterbank-to-DSU and 2) DSU-to-Translation data, and take the encoder from 1) and the decoder from 2) to initialise a new model, finetuning this on limited speech-translation data. The final model becomes compact by using the DSU pretraining to distil the knowledge of the SSL model. Our method has several benefits over using DSU as model inputs, such as shorter inference pipeline and robustness over (DSU) tokenization. In contrast to ASR pretraining, it does not require transcripts, making it applicable to low-resource settings. Evaluation on CoVoST-2 X-En shows that our method is >$0.5$ BLEU better than a ST model that directly finetune the SSL model, given only half the model size, and on a par with ASR pretraining.
Compact Speech Translation Models via Discrete Speech Units Pretraining
Tsz Kin Lam and Alexandra Birch and Barry Haddow School of Informatics, University of Edinburgh {tlam, a.birch, bhaddow}@ed.ac.uk
1 Introduction
In Speech-to-Text Translation (ST), the use of SSL models, e.g., wav2vec2 and HuBERT Baevski et al. (2020); Hsu et al. (2021), as model initialization is now common to obtain SOTA result Agarwal et al. (2023). Nevertheless, using these bulky models as initialisation imposes a large memory footprint to the downstream ST, which limits flexibility in modelling choices. In addition, it hinders on-device deployment that is crucial for privacy and useful in the absence of internet connection.
Knowledge distillation (KD) Kim and Rush (2016); Inaguma et al. (2021) is the standard approach to creating smaller models, but obtaining the high-quality pseudo labels for training ST systems is very costly. Instead, we focus on Discrete Speech Units (DSU) since they can be obtained cheaper, e.g., without requiring a translation model, and are an intermediate representation between speech and text. DSUs are K-means clusters of speech representations from selected layers within a SSL model. It represents sequences of discrete tokens, which are easier to model within a text processing architecture Polyak et al. (2021); Chou et al. (2023). DSU sequences are far smaller than dense Filterbank sequences, and they can be shortened with removing duplicates and Byte Pair Encoding (BPE) Sennrich et al. (2016). Therefore, in ST, a straightforward method to distill the SSL models is to replace the dense Filterbank by the DSUs, aka the DSU-to-Translation (DSU-to-Trl) model. Although using DSUs for training allows for transfer learning, using them at inference requires large and slow models.
To address the above, we use DSU for pretraining rather than as model input to make ST models more compact. Our method leverages large SSL models by pretraining smaller models on the corresponding DSU. More specifically, it pretrains encoder-decoder models on 1) Filterbank-to-DSU (Fbk-to-DSU) and 2) DSU-to-Trl data, and takes the encoder from 1) and the decoder from 2) to initialise a new model. After that, the new model is finetuned on limited speech-translation data. Under this formulation, (1) the model is much more compact than the SSL model (2) the compact model does not use DSU as inputs which (a) avoids the lengthy SSL and clustering inference pipeline and (b) minimizes its sensitivity to (DSU) tokenization. (3) Our method does not require transcripts, unlike ASR pretraining, making it applicable to low-resource settings. Similar to ST methods that use pretrained components, our method could be limited by the pretraining modality gap Liu et al. (2020); Le et al. (2023). Motivated by prior works, we investigate mitigating it with Connectionist Temporal Classification (CTC) Graves et al. (2006) regularisation.
We evaluate our method on CoVoST-2 Wang et al. (2021) X-En (21 language directions) with multilingual ST. By simply using a monolingual HuBERT model to extract the DSU, our method has a gain of >$3$ BLEU with respect to training from scratch, demonstrating DSU’s cross-lingual ability. Our method is also >$0.5$ BLEU better than a ST model that directly finetuned HuBERT despite having half model size. Our main contributions are:
  • We propose to distill knowledge from large SSL models to make ST models more compact via DSU pretraining. Our method avoids the lengthy inference pipeline in the DSU-to-Trl method.
  • We examine the effect of (DSU) tokenization on the ST models. Our method is robust across different tokenizations, which is not the case for the DSU-to-Trl method.
  • We examine applying CTC to the DSU in pretraining and on the translation in finetuning to mitigate the modality gap, and we show that both improve performance.
2 Related Work
There are a number of related works that use DSU to enhance ST. Fang and Feng (2023) and Zhang et al. (2023b) use DSU to create more training data in a back-translation fashion. Chang et al. (2023) and Zhang et al. (2023b) explore the replacement of Filterbank by DSU as speech input. In contrast, we use DSU and its translation model for pretraining, resulting in a better Fbk-to-Trl model that has a shorter inference pipeline.
In the case of pretraining, Wu et al. (2023) use a single Speech-to-DSU model in pretraining for general speech-to-text purposes whereas we tailor the use for ST by using a pair of encoder-decoder models. Zhang et al. (2022b) also decompose ST into speech-to-unit and unit-to-text tasks. Their training is based on masked unit prediction, and it requires an extra unit-encoder module in inference. In contrast, we resort to supervised training on the DSU in acoustic pretraining and require no extra module in inference. More importantly, our goal is to make (multilingual) ST more compact, aiming also at low-resource settings where transcripts are not easily available, rather than learning a joint semantic space for both transcripts and audios.
AST model (#Params) High Mid Low All
Scratch (52M) $19.37$ $7.91$ $0.73$ $5.99$
ASR-Pretraining (52M) $26.53$ $12.22$ $1.82$ $9.00$
Hu-Transformer (113M) $24.31$ $11.42$ $2.18$ $8.60$
DSU-Adapter (48M) $26.54$ $12.91$ $1.76$ $9.13$
Table 1: Results in BLEU on CoVoST-2 X-En. DSU-Adapter is 0.53 points better than Hu-Transformer that is $2.3$ times larger and on a par with ASR-Pretraining.
3 Method
Our method distills knowledge from the SSL representations via first DSU pretraining (PT) and then finetuning (FT) on the paired ST data.
In PT, our method uses DSU from an SSL speech model to train a pair of encoder-decoder models: 1) Fbk-to-DSU and 2) DSU-to-Trl. The Fbk-to-DSU model takes the log Mel Filterbank as the encoder input and predicts the corresponding DSU sequence. The model is trained by Maximum Likelihood Estimation (MLE). In addition, we interpolate between MLE and CTC applied to the last encoder layer:
$\mathcal{L}^{\text{Fbk-to-DSU}}=(1-\lambda_{\alpha})\mathcal{L}_{\text{MLE}}(\mathbf{U}|\mathbf{F})+\lambda_{\alpha}\mathcal{L}_{\text{CTC}}(\mathbf{\tilde{U}}|\mathbf{F})$ (1)
[…]
In FT, we use the encoder of the Fbk-to-DSU model and the decoder (and its output layer) of the DSU-to-Trl model to initialise the compact model, followed by finetuning on the paired speech-translation data using both MLE and CTC loss Gaido et al. (2021); Zhang et al. (2023a):
$\mathcal{L}^{\text{FT}}=(1-\lambda_{\beta})\mathcal{L}_{\text{MLE}}(\mathbf{Y}|\mathbf{F})+\lambda_{\beta}\mathcal{L}_{\text{CTC}}(\mathbf{\tilde{Y}}|\mathbf{F})$ (2)
[…]
Tokenization
As mentioned in section 1, we examine the tokenization effect, including using separate (source/target) vocabularies or not. We denote each DSU token by a hashtag with an index, e.g,. ’#456’. Following Ao et al. (2022), we reduce the sequence length via merging the sequential repetitions, e.g, ’#1 #1 #1 #456 #456 #23’ becomes ’#1 #456 #23’. When BPE is applied on the DSU, the whitespace characters are ignored, e.g., ’#1 #456 #23’ becomes ’#1#456#23’ which could be split into ’#1#45 6#2 3’.
Has BPE $|\mathcal{V}|$ DSU Length DSU-to-Trl Enc-Init EncDec-Init DSU-Adapter
on DSU? Length Ratio $20$M to $27$M $52$M to $70$M $46$M to $64$M $48$M to $67$M
No $1$K-$8$K $176$ $12.9$ $6.73$ $7.70$ $7.87$ $8.54$
$1$K-$16$K " $14.1$ $6.36$ $7.50$ $8.00$ $8.43$
$1$K-$32$K " $14.9$ $6.30$ $7.23$ $7.65$ $7.94$
$8$K " $12.7$ $6.88$ $7.64$ $8.05$ $8.26$
$16$K " $14.0$ $6.33$ $7.41$ $7.91$ $8.17$
$32$K " $14.9$ $6.26$ $6.66$ $7.41$ $7.68$
Yes $1$K-$8$K $221$ $16.3$ $4.52$ $8.23$ $8.44$ $8.61$
$16$K-$8$K $129$ $9.5$ $5.06$ $8.51$ $8.76$ $8.95$
$32$K-$8$K $115$ $8.5$ $4.43$ $8.67$ $9.02$ $9.13$
$8$K $150$ $7.6$ $7.02$ $8.33$ $8.51$ $8.82$
$16$K $133$ $7.7$ $6.50$ $8.57$ $8.61$ $8.93$
$32$K $118$ $7.8$ $5.07$ $8.30$ $8.44$ $8.70$
Table 2: Tokenization effect on 4 types of ST methods. In general, darker (brighter) cells refer to weaker (stronger) models. The best two models apply both BPE on the DSU and separate vocabularies in PT (cells in yellow).
4 Experiments
4.1 Data Preprocessing
We follow standard practices to preprocess the CoVoST-2 X-En data, i.e., log-Mel Filterbank of $80$ dimensions for speech inputs and BPE with $8$K vocabulary on the English target, $16$K on the transcripts and $32$K on the DSU, unless otherwise specified. We use an English HuBERT-Base model to extract the DSU by first downsampling the CoVoST-2 audio to $16$KHz. Each audio data utterance is then converted into the DSU, i.e., the clustering indexes, by applying K-Means clustering (K=$1,000$) on its HuBERT representation from the 6th layer Lakhotia et al. (2021). We sample training data from each language pair to form a subset of $98$K instances to train the K-Means model. Further details are in Appendix A.1.
4.2 Model Configuration
All models are based on Transformer Vaswani et al. (2017) with implementations from fairseq Ott et al. (2019); Wang et al. (2020). In the Fbk-to-token [^1] Token refers to either transcriptions, DSU, or translations. models, the encoder has convolutional layers to downsample the Fbk by a factor of 4. There are 12-6 layers in the transformer encoder-decoder, whereas the embedding and feed-forward network (FFN) dimensions are 256 and 4,096 respectively, unless otherwise specified. It is worth noting that: (1) The Fbk-to-DSU model is not trained on the translations, so it is not directly comparable to the ST models. Its effect to ST lies on its pretrained encoder (Table 2). (2) The DSU-to-Trl model is a ST model, and it can also be used for initialization.
Scratch is a ST model trained on the paired speech-translation data without pretraining.
ASR Pretraining refers to a ST model whose encoder is initialized by a speech recognition task with CTC regularisation on the transcripts.
DSU-to-Trl follows the Transformer used in text translation. We use 6-6 layers in the encoder-decoder which the dimension of embedding and FFN is 256 and 2,048 respectively. Despite its smaller model size, its inference requires storing and calling both HuBERT and K-Means models.
Hu-Transformer uses the entire HuBERT as the speech encoder initialization Fang and Feng (2023). For comparison to our DSU-Adapter, its subsequent encoder-decoder also has 1-6 layers.
DSU-Adapter is our proposed method. To better align the two pre-trained components, we also experiment with adding an extra encoder layer [^2] All parameters are trainable in ST finetuning. after the pre-trained encoder. Since its decoder is initialized by the DSU-to-Trl method, its decoder FFN dimension is $2,048$.
Enc-Init is a ST model that has its encoder initialized by the Fbk-to-DSU encoder. EncDec-Init is a DSU-Adapter model without the adapter layer.
4.3 Training and Inference
We leave the details in Appendix A.3. It is worth noting that we do not use extra audio data, e.g., Libri-Light Kahn et al. (2020) in pretraining.
5 Results and Analysis
We divide the 21 language pairs by resource level into: 1) "High": {ca, de, es, fr}, 2) "Mid": {fa, it, pt, ru and zh}, 3) "Low": {ar, cy, et, id, ja, lv, mn, nl, sl, sv, ta, tr} and 4)"All": the 21 languages pairs. We report the average BLEU over the test sets of each group using sacrebleu Post (2018)333nrefs:1—case:mixed—eff:no—tok:13a—smooth:exp—version:2.3.1.
Table 1 compares our DSU-Adapter and the baselines. Our DSU-Adapter is $3$ BLEU points (in the group "All") higher than Scratch. Furthermore, it is better than Hu-Transformer in spite of having half the parameters. For "Mid" and "High", the improvement is $1.49$ and $2.23$ points respectively, but it falls short by $0.42$ points for "Low". We also compare to ASR pretraining, which is not always applicable, e.g., in low-resource setting or perhaps even in an unwritten language Zhang et al. (2022a). Surprisingly, our adapter is on a par with it, i.e., $0.13$ points better. The result remains consistent when measured in chrF Popović (2015) and COMET [^4] WMT22-COMET-DA Rei et al. (2022), see Appendix A.4.
5.1 Tokenization and its Transfer
Table 2 presents the effect of tokenization. The 1st column indicates if BPE is applied on the DSU, and the 2nd column "$|\mathcal{V}|$" shows the vocabulary size, and if the model has a joint vocabulary or not in PT. For example, "$1$K-$8$K" means that we use a vocabulary of size $1$K for DSU and $8$K for English. These results are for all languages, please see Appendix A.4 for fine grained results.
Effect of BPE on DSU-to-Trl
When BPE is not applied on the DSU, the DSU-to-Trl models have a BLEU of $6.48$$\pm$ $0.26$ across the 6 vocabulary settings. Despite being smaller (<$30$M), these models are better than the Scratch model of $5.99$ BLEU.
When BPE is applied, the DSU sequence length becomes a crucial factor. On one hand, it delivers the best model ("$8$K": $7.02$ BLEU). On the other hand, the result deteriorates a lot in the other 5 settings, i.e., a BLEU of $5.12$$\pm$ $0.83$. More importantly, BPE’s correlation with the model performance is not straightforward, neither is the length ratio. For an example, the "$32$K" model (DSU length of $118$) is about $2$ BLEU points behind to the "$8$K" model (DSU length of $150$).
DSU-Adapter is more tokenization robust
When BPE is applied on the DSU, DSU-Adapter has a higher BLEU of $8.86$$\pm$ $0.19$, in contrast to $8.17$$\pm$ $0.32$. This behaviour is opposed to the DSU-to-Trl model that has a lower BLEU of $5.54$$\pm$ $1.07$ (with also larger variance), in contrast to $6.48$$\pm$ $0.26$ when BPE is not used. The better scores with smaller variance show that DSU-Adapter is more tokenization robust. We see this as a benefit of introducing the DSU (SSL knowledge) via PT rather than as model inputs. Besides, using separate vocabularies in PT is preferred since it performs slightly better, and the DSU, which are not needed in the ST output, would not occupy the target vocabulary.
Ablation
Having similar size, DSU-Adapter is better than EncDec-Init and Enc-Init. The BLEU over their 12 settings is $8.51$$\pm$ $0.44$, $8.22$$\pm$ $0.51$, and $7.99$$\pm$ $0.61$ respectively. Encoder-initialization seems more crucial than decoder-initialization, as reflected by the fact that the best DSU-Adapter comes from a combination with the weakest DSU-to-Trl model of $4.43$ BLEU.
5.2 CTC regularisation helps
Has CTC in High Mid Low All
DSU PT? ST FT?
No No $25.81$ $9.91$ $1.46$ $8.14$
No Yes $26.10$ $11.35$ $1.69$ $8.71$
Yes No $25.94$ $10.82$ $1.51$ $8.44$
Yes Yes $26.12$ $11.53$ $1.73$ $8.74$
Table 3: Effect of CTC on Fbk-to-DSU PT and/or ST FT to the DSU-Adapter method.
Table 3555Owing to the large number of configurations, we can experiment on 4 vocabularies only and then report their average. presents an analysis of using CTC in the Fbk-to-DSU PT and/or in the ST FT using the DSU-Adapter method. We find that CTC helps on either stage, but the gain is 0.27 BLEU more in finetuning. Using them jointly still helps, but the gain becomes marginal, i.e, 0.03 BLEU.
6 Conclusion
We propose to distill knowledge from the large SSL model via pretraining on its DSU to make ST models more compact for on-device deployment. Our compact model, i.e., the DSU-Adapter, shows stronger result over finetuning the corresponding (larger) HuBERT in multilingual ST task. In contrast to using DSU as model input, our method requires less components in inference and shows stronger robustness over the tokenization. Finally, our method is suitable for the low-resource setting since it requires no transcripts.
7 Limitations
In this work, we propose to distill knowledge from the informative SSL representations via DSU pretraining to make ST models more compact. In addition, we also discuss advantages of our method over using DSU as model inputs. In spite of its noticeable benefits, there are several factors that are not thoroughly explored:
K-Means Clustering
We did not inspect the clustering size (fixed to $1,000$) and the number of training instances (only fixed to $98,000$) used in training the K-Means clustering model.
Other acoustic encoders
We did not experiment other acoustic encoders, such as conformer Gulati et al. (2020); Papi et al. (2023) and E-Branchformer Peng et al. (2023). This stronger encoders should provide further gains for our method since they also enjoy the benefit of pretraining.
Self-Supervised Models
We did not experiment using other layers (only fixed to the 6th layer) from the HuBERT model nor from stronger SSL models to extract the DSU. These two factors potentially increase the DSU quality and potentially provide better improvement for our method while maintaining the same model size.
Acknowledgements
This work was funded by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee (grant number 10039436: UTTER). The computations described in this research were performed using the Baskerville Tier 2 HPC service (https://www.baskerville.ac.uk/). Baskerville was funded by the EPSRC and UKRI through the World Class Labs scheme (EP/T022221/1) and the Digital Research Infrastructure programme (EP/W032244/1) and is operated by Advanced Research Computing at the University of Birmingham.
Appendix A Appendix
A.1 Data preprocessing
log Mel Filterbank
The 80-D log Mel Filterbank is computed every 10ms with a 25ms window, followed by feature normalisation using statistics, i.e., mean and variance, computed over each channel.
BPE
We use the BPE implementation from sentencepiece Kudo and Richardson (2018) and obtain vocabulary of size 8K on the English target, 16K on the transcripts and 32K on the DSU, unless otherwise specified.
HuBERT
K-Means Clustering
We use MiniBatchKMeans from sklearn with a clustering size of $1,000$ and default settings in other hyper-parameters. When training the clustering model, we divide the 21 language pairs into three groups: 1) {ar, cy, et, id, ja, lv, mn, sl, sv, ta, tr}, 2) {nl, pt, ru, zh} and 3) {ca, de, es, fa, fr, it} for sampling. For each language pair in group 1), we sample $1$K instances, whereas it becomes $3$K in group 2) and $12.5$K in group 3), resulting in a total of $98$K instances.
A.2 Model Configuration
We use "pre" layer-normalization Nguyen and Salazar (2019) in DSU-to-Trl.
A.3 Training and Inference
In training,
  • we skip training data that are more than $3$K frames for Filterbank, $320$K frames for HuBERT input or $1,024$ tokens.
  • we apply SpecAugment Park et al. (2019) with parameters: {$F=30,T=40,m_{F}=2,m_{T}=2$} on Filterbank inputs.
  • we share the embedding weights when using a joint vocabulary in the DSU-to-Trl model.
  • we set $\lambda_{\alpha}$ and $\lambda_{\beta}$ in CTC to $0.3$ and the smoothing parameter in LS to $0.1$
  • we initialize the encoder (decoder) with the last (best) checkpoint from the PT model.
All models are optimized by Adam Kingma and Ba (2015) with inverse square root scheduler. For simplicity, we provide major training parameters for each model in the format of {effective batch size, warm-up steps, peak learning rate, maximum training steps}:
Fbk-to-Token:
{32K frames, 25K, 2e-3, 60K}
DSU-to-Trl:
{80K tokens, 10K, 5e-4, 50K}
Hu-Transformer:
{4M frames, 4K, 1e-4, 300K}
In inference, we average the last 5 checkpoints and use beam size of 5 in generation. All experiments are run on Nvidia A100 GPUs. It takes about 1 day for 2 A100 (40GB) GPUs to complete an experiment that uses Filterbank as inputs.
A.4 Results and Analysis
In addition to BLEU (see Table 8 for the result of each language), we also provide the result of Table 1 using chrF777nrefs:1—case:mixed—eff:yes—nc:6—nw:0—space:no—version:2.3.1 and COMET. In order to compute the COMET score, we feed the gold-reference transcripts as the source sentence.
Table 4 and Table 6 show the average result in chrF and COMET respectively over the four groups defined in section 5. Their results follow the findings in BLEU: the DSU-Adapter is better than Hu-Transformer and on a par with ASR pretraining. Further details, such as training data size (in terms of the number of utterances rather than hours888”zh” has 10 hours of audio data, whereas ”nl” has only 7 hours. Therefore, ”zh” is classified into ”Mid” resource group even though it has slightly less number of utterances than ”nl”.) and per language performance, are shown on Table 5 for chrF and Table 7 for COMET.
CTC Regularisation
On Table 3, we average the results over this four configurations: 1) "No-BPE 1K-8K", 2) "BPE 8K", 3) "BPE 32K" and 4) "BPE 32K-8K".
AST model Transcripts Model size High Mid Low All
Scratch No 52M $43.58$ $27.22$ $14.62$ $23.10$
ASR Pretraining Yes 52M $51.90$ $32.82$ $16.38$ $27.10$
Hu-Transformer No 113M $49.88$ $31.94$ $16.96$ $26.80$
DSU-Adapter No 48M $52.08$ $33.94$ $16.46$ $27.40$
Table 4: Results of Table 1 in chrF score by resource group. Similar to the findings in BLEU, DSU-Adapter is far better than Hu-Transformer, in particular for "High" and "Mid", and on a par with ASR Pretraining. The best result in each group is denoted by ’_’.
Languages #Training data Resources Scratch ASR-PT Hu-Transformer DSU-Adapter
ca 95,854 High $42.5$ $50.9$ $49.8$ $51.5$
de 127,824 $35.4$ $45.1$ $43.6$ $45.4$
es 79,013 $48.1$ $56$ $52.8$ $55.7$
fr 207,372 $48.3$ $55.6$ $53.3$ $55.7$
fa 53,949 Mid $15.6$ $16.9$ $19.9$ $16.9$
it 31,698 $40.6$ $50.2$ $46.4$ $50.3$
pt 9,158 $25$ $31.3$ $33.8$ $32.3$
ru 12,112 $35.6$ $42.6$ $40$ $47.7$
zh 7,085 $19.3$ $23.1$ $19.6$ $22.5$
ar 2,283 Low $12.5$ $14.4$ $13.8$ $13.2$
cy 1,241 $14.5$ $16.3$ $15.8$ $15.4$
et 1,782 $17.9$ $21.3$ $19$ $19.9$
id 1,243 $13.8$ $14.2$ $14.9$ $14.5$
ja 1,119 $14$ $14$ $13.7$ $14.4$
lv 2,337 $13.9$ $14.5$ $15.9$ $15.4$
mn 2,067 $14.5$ $15.4$ $14.3$ $14.8$
nl 7,108 $17.8$ $24.3$ $30.8$ $27$
sl 1,843 $13.9$ $15.4$ $15.3$ $15.4$
sv 2,160 $13.2$ $13.8$ $17.3$ $14.5$
ta 1,358 $12$ $11.9$ $12$ $12.3$
tr 3,966 $17.4$ $21.1$ $20.7$ $20.7$
Table 5: Results of Table 1 in chrF score for each language pair. The best result in each group is denoted by ’_’.
AST model Transcripts Model size High Mid Low All
Scratch No 52M $0.605$ $0.498$ $0.433$ $0.481$
ASR Pretraining Yes 52M $0.680$ $0.537$ $0.443$ $0.511$
Hu-Transformer No 113M $0.650$ $0.522$ $0.439$ $0.499$
DSU-Adapter No 48M $0.681$ $0.548$ $0.442$ $0.513$
Table 6: Results of Table 1 in COMET score by resource group. Similar to the findings in BLEU, DSU-Adapter is far better than Hu-Transformer, in particular for "High" and "Mid", and on a par with ASR Pretraining. The best result in each group is denoted by ’_’.
Languages #Training data Resources Scratch ASR-PT Hu-Transformer DSU-Adapter
ca 95,854 High $0.607$ $0.681$ $0.654$ $0.684$
de 127,824 $0.535$ $0.608$ $0.588$ $0.613$
es 79,013 $0.620$ $0.696$ $0.654$ $0.692$
fr 207,372 $0.660$ $0.736$ $0.704$ $0.735$
fa 53,949 Mid $0.479$ $0.494$ $0.501$ $0.493$
it 31,698 $0.557$ $0.640$ $0.594$ $0.641$
pt 9,158 $0.468$ $0.497$ $0.510$ $0.508$
ru 12,112 $0.551$ $0.580$ $0.563$ $0.630$
zh 7,085 $0.434$ $0.472$ $0.440$ $0.467$
ar 2,283 Low $0.438$ $0.465$ $0.45$ $0.445$
cy 1,241 $0.420$ $0.448$ $0.424$ $0.426$
et 1,782 $0.413$ $0.432$ $0.407$ $0.423$
id 1,243 $0.420$ $0.406$ $0.416$ $0.412$
ja 1,119 $0.417$ $0.412$ $0.411$ $0.422$
lv 2,337 $0.457$ $0.453$ $0.465$ $0.461$
mn 2,067 $0.429$ $0.441$ $0.419$ $0.437$
nl 7,108 $0.427$ $0.460$ $0.483$ 0.477
sl 1,843 $0.430$ $0.428$ $0.433$ $0.428$
sv 2,160 $0.454$ $0.448$ $0.472$ $0.454$
ta 1,358 $0.413$ $0.406$ $0.403$ $0.411$
tr 3,966 $0.480$ $0.519$ $0.484$ $0.514$
Table 7: Results of Table 1 in COMET score for each language pair. The best result in each group is denoted by ’_’.
Languages #Training data Resources Scratch ASR-PT Hu-Transformer DSU-Adapter
ca 95,854 High $18.24$ $25.66$ $24.24$ $25.98$
de 127,824 $13.97$ $21.06$ $19.28$ $21.32$
es 79,013 $22.37$ $29.8$ $26.64$ $29.34$
fr 207,372 $22.91$ $29.6$ $27.09$ $29.52$
fa 53,949 Mid $1.9$ $3.07$ $4.63$ $2.84$
it 31,698 $15.1$ $23.55$ $19.98$ $23.36$
pt 9,158 $7.01$ $11.2$ $13.13$ $11.69$
ru 12,112 $11.97$ $16.87$ $14.95$ $21$
zh 7,085 $3.58$ $6.39$ $4.4$ $5.67$
ar 2,283 Low $0.44$ $1.96$ $1.26$ $0.48$
cy 1,241 $0.64$ $1.24$ $1.21$ $0.72$
et 1,782 $0.49$ $1.46$ $0.61$ $0.93$
id 1,243 $0.31$ $0.67$ $1.27$ $0.79$
ja 1,119 $0.27$ $0.67$ $0.36$ $0.62$
lv 2,337 $1.22$ $1.65$ $2.63$ $2.06$
mn 2,067 $0.16$ $0.44$ $0.31$ $0.19$
nl 7,108 $2.14$ $6.95$ $10.75$ $8.46$
sl 1,843 $0.41$ $1.83$ $1.23$ $1.81$
sv 2,160 $0.9$ $1.28$ $3.4$ $1.76$
ta 1,358 $0.14$ $0.14$ $0.13$ $0.12$
tr 3,966 $1.67$ $3.5$ $3.04$ $3.14$
Table 8: Results of Table 1 in BLEU score for each language pair. The best result in each group is denoted by ’_’.

### v2

@@
 Abstract
-Using Self-Supervised Learning (SSL) as model initialization is now common to obtain strong results in Speech Translation (ST).
-However, they also impose a large memory footprint, hindering on-device deployment.
-In this paper, we leverage the SSL models by pretraining smaller models on their Discrete Speech Units (DSU).
-We pretrain encoder-decoder models on 1) Filterbank-to-DSU and 2) DSU-to-Translation data, and take the encoder from 1) and the decoder from 2) to initialise a new model, finetuning this on limited speech-translation data.
-The final model becomes compact by using the DSU pretraining to distil the knowledge of the SSL model.
-Our method has several benefits over using DSU as model inputs, such as shorter inference pipeline and robustness over (DSU) tokenization.
-In contrast to ASR pretraining, it does not require transcripts, making it applicable to low-resource settings.
-Evaluation on CoVoST-2 X-En shows that our method is >$0.5$ BLEU better than a ST model that directly finetune the SSL model, given only half the model size, and on a par with ASR pretraining.
+We propose a pretraining method to use Self-Supervised Speech (SSS) model to creating more compact Speech-to-text Translation.
+In contrast to using the SSS model for initialization, our method is more suitable to memory constrained scenario such as on-device deployment.
+Our method is based on Discrete Speech Units (DSU) extracted from the SSS model.
+In the first step, our method pretrains two smaller encoder-decoder models on 1) Filterbank-to-DSU (Fbk-to-DSU) and 2) DSU-to-Translation (DSU-to-Trl) data respectively.
+The DSU thus become the distillation inputs of the smaller models.
+Subsequently, the encoder from the Fbk-to-DSU model and the decoder from the DSU-to-Trl model are taken to initialise the compact model.
+Finally, the compact model is finetuned on the paired Fbk-Trl data.
+In addition to being compact, our method requires no transcripts, making it applicable to low-resource settings.
+It also avoids speech discretization in inference and is more robust to the DSU tokenization.
+Evaluation on CoVoST-2 (X-En) shows that our method has consistent improvement over the baseline in three metrics while being compact i.e., only half the SSS model size.
 Compact Speech Translation Models via Discrete Speech Units Pretraining
@@
 1 Introduction
-In Speech-to-Text Translation (ST), the use of SSL models, e.g., wav2vec2 and HuBERT Baevski et al.
+In Speech-to-text Translation (ST), using Self-Supervised Speech (SSS) models, such as wav2vec 2.0 and HuBERT Baevski et al.
 (2020); Hsu et al.
-(2021), as model initialization is now common to obtain SOTA result Agarwal et al.
+(2021), as model initialization is now common to obtain the SOTA result Agarwal et al.
 (2023).
-Nevertheless, using these bulky models as initialisation imposes a large memory footprint to the downstream ST, which limits flexibility in modelling choices.
-In addition, it hinders on-device deployment that is crucial for privacy and useful in the absence of internet connection.
-Knowledge distillation (KD) Kim and Rush (2016); Inaguma et al.
-(2021) is the standard approach to creating smaller models, but obtaining the high-quality pseudo labels for training ST systems is very costly.
-Instead, we focus on Discrete Speech Units (DSU) since they can be obtained cheaper, e.g., without requiring a translation model, and are an intermediate representation between speech and text.
-DSUs are K-means clusters of speech representations from selected layers within a SSL model.
-It represents sequences of discrete tokens, which are easier to model within a text processing architecture Polyak et al.
+Nevertheless, such model initialisation makes the ST model less memory-adaptive and could impose a large memory footprint.
+These factors hinders on-device deployment that is crucial for privacy and useful in the absence of internet connection.
+How can we use the SSS model(s) to create a more compact ST model?
+When using the SSS model for initialization, the corresponding ST model uses the dense representations of the SSS model for its task.
+Alternatively, an informative proxy, which requires less memory to obtain, for the dense representation may make the ST model more compact.
+Discrete Speech Units (DSU) extracted from the SSS model can be such a good proxy.
+DSU are K-Means clusters of speech representations from selected layers of the SSS model.
+It represents sequence of discrete tokens, which are easier to model within a text processing architecture Polyak et al.
 (2021); Chou et al.
 (2023).
-DSU sequences are far smaller than dense Filterbank sequences, and they can be shortened with removing duplicates and Byte Pair Encoding (BPE) Sennrich et al.
-(2016).
-Therefore, in ST, a straightforward method to distill the SSL models is to replace the dense Filterbank by the DSUs, aka the DSU-to-Translation (DSU-to-Trl) model.
-Although using DSUs for training allows for transfer learning, using them at inference requires large and slow models.
-To address the above, we use DSU for pretraining rather than as model input to make ST models more compact.
-Our method leverages large SSL models by pretraining smaller models on the corresponding DSU.
-More specifically, it pretrains encoder-decoder models on 1) Filterbank-to-DSU (Fbk-to-DSU) and 2) DSU-to-Trl data, and takes the encoder from 1) and the decoder from 2) to initialise a new model.
-After that, the new model is finetuned on limited speech-translation data.
-Under this formulation, (1) the model is much more compact than the SSL model (2) the compact model does not use DSU as inputs which (a) avoids the lengthy SSL and clustering inference pipeline and (b) minimizes its sensitivity to (DSU) tokenization.
-(3) Our method does not require transcripts, unlike ASR pretraining, making it applicable to low-resource settings.
-Similar to ST methods that use pretrained components, our method could be limited by the pretraining modality gap Liu et al.
-(2020); Le et al.
-(2023).
-Motivated by prior works, we investigate mitigating it with Connectionist Temporal Classification (CTC) Graves et al.
-(2006) regularisation.
+DSU sequences [^1] In this paper, DSU and DSU sequences are used interchangeably.
+When we need to focus on a few units of the sequence, we call them DSU tokens.
+are far smaller than the sequences of dense representations.
+Therefore, a straightforward method to distill the SSS models is to use DSU as speech inputs, aka the DSU-to-Translation (DSU-to-Trl) model.
+Although using DSU as inputs allows for transfer learning and a memory-adaptive model, using them at inference still requires storing and calling the quantization modules, i.e, the SSS model and the K-Means model.
+We thus propose to use DSU for pretraining (PT) rather than as model input to make ST models more compact.
+Our method distils the SSS model by pretraining smaller models on the corresponding DSU.
+More specifically, our method firstly pretrains two smaller encoder-decoder models on 1) Filterbank-to-DSU (Fbk-to-DSU) and 2) DSU-to-Trl data respectively.
+The DSU thus become the distillation inputs of the smaller models.
+Subsequently, the encoder from the Fbk-to-DSU model and the decoder from the DSU-to-Trl model are taken to initialise the compact model.
+Finally, the compact model is finetuned on the paired Fbk-Trl data.
+Under this formulation, (1) we can use the SSS model to create a ST model that is adaptive to the memory footprint.
+(2) Our method requires no transcripts, unlike ASR-pretraining, making it applicable to low-resource settings.
+(3) Our method avoids using the quantization modules in inference.
+(4) Extensive results also show that our method is more robust to DSU tokenization than the DSU-to-Trl method.
 We evaluate our method on CoVoST-2 Wang et al.
-(2021) X-En (21 language directions) with multilingual ST.
-By simply using a monolingual HuBERT model to extract the DSU, our method has a gain of >$3$ BLEU with respect to training from scratch, demonstrating DSU’s cross-lingual ability.
-Our method is also >$0.5$ BLEU better than a ST model that directly finetuned HuBERT despite having half model size.
+(2021) X-En language directions ($21$ in total) using multilingual ST.
+By using a HuBERT-Base model to extract the DSU, our method shows strong and consistent improvements in three evaluation metrics with respect to a ST model that is trained from scratch.
 Our main contributions are:
-  • We propose to distill knowledge from large SSL models to make ST models more compact via DSU pretraining. Our method avoids the lengthy inference pipeline in the DSU-to-Trl method.
-  • We examine the effect of (DSU) tokenization on the ST models. Our method is robust across different tokenizations, which is not the case for the DSU-to-Trl method.
-  • We examine applying CTC to the DSU in pretraining and on the translation in finetuning to mitigate the modality gap, and we show that both improve performance.
+  • We propose a pretraining method to distil the SSS model to creating a more compact ST model. Rather than competing with the SOTA ST models, adaptability to the memory footprint is our key focus.
+  • Our method uses DSU for pretraining rather than as model inputs. This lowers the inference cost, especially for on-device purpose, by avoiding the quantization modules (storage and running).
+  • We conduct extensive analysis to study the effect of DSU tokenization to both using DSU as model inputs and as pretraining.
+Our pretraining method is found to be more robust to different tokenizations.
 2 Related Work
@@
 (2023b) explore the replacement of Filterbank by DSU as speech input.
+Furthermore, Yan et al.
+(2024) proposes a multi-tasking learning framework with hard parameter sharing, i.e., using a joint vocabulary for text tokens and DSU, to improve the speech-text modality gap.
 In contrast, we use DSU and its translation model for pretraining, resulting in a better Fbk-to-Trl model that has a shorter inference pipeline.
@@
 More importantly, our goal is to make (multilingual) ST more compact, aiming also at low-resource settings where transcripts are not easily available, rather than learning a joint semantic space for both transcripts and audios.
-AST model (#Params) High Mid Low All
-Scratch (52M) $19.37$ $7.91$ $0.73$ $5.99$
-ASR-Pretraining (52M) $26.53$ $12.22$ $1.82$ $9.00$
-Hu-Transformer (113M) $24.31$ $11.42$ $2.18$ $8.60$
-DSU-Adapter (48M) $26.54$ $12.91$ $1.76$ $9.13$
-Table 1: Results in BLEU on CoVoST-2 X-En. DSU-Adapter is 0.53 points better than Hu-Transformer that is $2.3$ times larger and on a par with ASR-Pretraining.
 3 Method
-Our method distills knowledge from the SSL representations via first DSU pretraining (PT) and then finetuning (FT) on the paired ST data.
-In PT, our method uses DSU from an SSL speech model to train a pair of encoder-decoder models: 1) Fbk-to-DSU and 2) DSU-to-Trl.
-The Fbk-to-DSU model takes the log Mel Filterbank as the encoder input and predicts the corresponding DSU sequence.
-The model is trained by Maximum Likelihood Estimation (MLE).
-In addition, we interpolate between MLE and CTC applied to the last encoder layer:
-$\mathcal{L}^{\text{Fbk-to-DSU}}=(1-\lambda_{\alpha})\mathcal{L}_{\text{MLE}}(\mathbf{U}|\mathbf{F})+\lambda_{\alpha}\mathcal{L}_{\text{CTC}}(\mathbf{\tilde{U}}|\mathbf{F})$ (1)
+Figure 1: Illustration of the Fbk-to-DSU model.
+It is like an auto-encoding training process, but between a continuous format (log Mel Filterbank) and its discrete format (DSU) that is extracted from a HuBERT model.
+Our method uses DSU in the form of pretraining to distil knowledge from the SSS (dense) representations to creating a more compact ST model.
+In the first step, our method pretrains two smaller encoder-decoder models on 1) Fbk-to-DSU and 2) DSU-to-Trl data respectively.
+The Fbk-to-DSU model takes the log Mel Fbk as the encoder input and predicts the DSU sequence.
+The model is trained by an interpolation of Connectionist Temporal Classification (CTC, Graves et al.
+(2006)) loss that is applied to the last encoder layer and label-smoothed Cross-Entropy (CE) loss:
+$\mathcal{L}^{\text{Fbk-to-DSU}}=(1-\lambda_{\alpha})\mathcal{L}_{\text{CE}}(\mathbf{U}|\mathbf{F})+\lambda_{\alpha}\mathcal{L}_{\text{CTC}}(\mathbf{\tilde{U}}|\mathbf{F})$ (1)
 […]
-In FT, we use the encoder of the Fbk-to-DSU model and the decoder (and its output layer) of the DSU-to-Trl model to initialise the compact model, followed by finetuning on the paired speech-translation data using both MLE and CTC loss Gaido et al.
+Next, we use the encoder of the Fbk-to-DSU model and the decoder (and its output layer) of the DSU-to-Trl model to initialise the compact model, followed by finetuning on the paired Fbk-Trl data using both CE and CTC loss Gaido et al.
 (2021); Zhang et al.
-(2023a):
-$\mathcal{L}^{\text{FT}}=(1-\lambda_{\beta})\mathcal{L}_{\text{MLE}}(\mathbf{Y}|\mathbf{F})+\lambda_{\beta}\mathcal{L}_{\text{CTC}}(\mathbf{\tilde{Y}}|\mathbf{F})$ (2)
+(2023a) on the translations:
+$\mathcal{L}^{\text{FT}}=(1-\lambda_{\beta})\mathcal{L}_{\text{CE}}(\mathbf{Y}|\mathbf{F})+\lambda_{\beta}\mathcal{L}_{\text{CTC}}(\mathbf{\tilde{Y}}|\mathbf{F})$ (2)
 […]
-Tokenization
-As mentioned in section 1, we examine the tokenization effect, including using separate (source/target) vocabularies or not.
-We denote each DSU token by a hashtag with an index, e.g,.
-’#456’.
-Following Ao et al.
-(2022), we reduce the sequence length via merging the sequential repetitions, e.g, ’#1 #1 #1 #456 #456 #23’ becomes ’#1 #456 #23’.
-When BPE is applied on the DSU, the whitespace characters are ignored, e.g., ’#1 #456 #23’ becomes ’#1#456#23’ which could be split into ’#1#45 6#2 3’.
+3.1 Tokenization of DSU in different models
+Figure 2: Aligning the DSU tokenization of the Fbk-to-DSU, DSU-to-Trl and compact ST model.
+The discrete nature of DSU makes the above training process similar to the transcripts-based pretraining.
+However, DSU is self-supervised, whereas transcripts require human annotations.
+DSU are also much longer and can be represented with various sets of symbols.
+The length issue could be relieved by merging sequential repetitions Ao et al.
+(2022), e.g., ’#$1$$1$$1$$456$$456$$23$’ becomes ’#$1$$456$$23$’, where each DSU token is denoted by a #{integer}.
+Byte Pair Encoding (BPE) Sennrich et al.
+(2016) could be applied to reduce the DSU sequence length further, e.g., ’#$1$$456$$23$’ could be split into a single subword unit: ’#$1$#$456$#$23$’.
+Since both Fbk-to-DSU and DSU-to-Trl models map to different targets, and DSU can be represented with various set of symbols, we align the tokenizations (or called vocabularies [^2] We use vocabulary and tokenization interchangeably, since we did not apply subword regularisation.) of the two models.
+Figure 2 provides an illustration.
+The vocabulary of the Fbk-to-DSU model (Vocabulary A) is identical to the source vocabulary of the DSU-to-Trl model (their weights are not shared since these two models are trained independently), whereas the target vocabulary (Vocabulary B) of the DSU-to-Trl model is identical to the target vocabulary of the final compact model (their weights are shared during initialisation).
+The DSU-to-Trl model is similar to a text translation model, so we also experiment of using separate vocabularies or a joint vocabulary.
+If a joint vocabulary of English subword units and DSU (BPE or not) is used, all the three models would have the same vocabulary, and the weights of the source and target vocabularies of the DSU-to-Trl model are also tied.
+4 Experiments
+4.1 Data Preprocessing
+We follow standard practices to preprocess the CoVoST-2 X-En data.
+For speech inputs using $80$-D log Mel Fbk, we computed the features for every $10$ms with a $25$ms window and then normalized them using its mean and variance computed over each channel.
+We use the BPE implementation from sentencepiece Kudo and Richardson (2018) and obtain vocabulary of size $8$K on the English target, $16$K on the (non-English) transcripts and $32$K on the DSU, unless otherwise specified.
+We use HuBERT-Base [^3] https://github.com/facebookresearch/fairseq/tree/main/examples/textless_nlp/gslm/speech2unit model to extract the DSU by first downsampling the CoVoST-2 audio to $16$KHz.
+Each audio data utterance is then converted into the DSU, i.e., the clustering indexes, by applying K-Means clustering (K=$1,000$; MiniBatchKMeans from sklearn) on its HuBERT representation from the $6$th layer Lakhotia et al.
+(2021).
+To train the K-Means model, we divide the $21$ language pairs into three groups: 1) {ar, cy, et, id, ja, lv, mn, sl, sv, ta, tr}, 2) {nl, pt, ru, zh} and 3) {ca, de, es, fa, fr, it}.
+We then sample $1$K instances for each language pair in group 1), which becomes $3$K in group 2) and $12.5$K in group 3), to create a multilingual training dataset of $98$K instances for the K-Means model.
+4.1.1 On the choice of using HuBERT-Base
+Given the rapid advance in the SSS models, there are many alternatives, such as XLS-R Babu et al.
+(2021) and Wavlm Chen et al.
+(2022), for extracting the DSU for our method.
+These models are larger in scale and could be multilingual, thus providing DSU of higher qualities.
+The improvement of our method by using DSU from the HuBERT-Base would probably be a lower-bound, considering its relatively poor qualities to the bigger models.
+Since our goal is about compactness via DSU pretraining rather than comparing the DSU qualities across the SSS models, we took a simple HuBERT-Base model to illustrate the idea.
+Pretraining only on English audio data could also suggest hints on whether the DSU and our method could be generalised to languages that are unseen to the SSS models.
+4.2 Model Configuration
+All models are based on Transformer Vaswani et al.
+(2017) with implementations from fairseq Ott et al.
+(2019); Wang et al.
+(2020).
+In the Fbk-to-Token (i.e.
+transcriptions, DSU, or translations) models, the encoder has convolutional layers to downsample the Fbk by a factor of $4$.
+There are $12$-$6$ layers in the transformer encoder-decoder, whereas the embedding and feed-forward network (FFN) dimensions are $256$ and $4,096$ respectively, unless otherwise specified.
+It is worth noting that: (1) The Fbk-to-DSU model is not trained on the translations, so it is not directly comparable to the ST models.
+Its effect on ST lies on its pretrained encoder (Table 2).
+(2) The DSU-to-Trl model is a ST model which decoder can be used for initialization.
+Scratch is a ST model trained on the paired speech-translation data without pretraining.
+ASR Pretraining refers to a ST model whose encoder is initialized by a speech recognition task with CTC regularisation on the transcripts.
+DSU-to-Trl follows the Transformer used in text translation.
+We use $6$-$6$ layers in the encoder-decoder which the dimension of embedding and FFN is $256$ and $2,048$ respectively.
+In addition, we use "pre" layer-normalization Nguyen and Salazar (2019).
+Despite its smaller model size, its inference requires the quantization modules.
+Hu-Transformer uses the entire HuBERT as the speech encoder initialization Fang and Feng (2023). For comparison to our DSU-Adapter, its subsequent encoder-decoder also has $1$-$6$ layers.
+DSU-Adapter is our proposed method.
+To better align the two pre-trained components, we also experiment with adding an extra encoder layer as a simple adapter layer after the pre-trained encoder.
+Because of the small model size, all model parameters are trainable.
+Since its decoder is initialized by the DSU-to-Trl method, its decoder FFN dimension is $2,048$.
+Enc-Init is a ST model that has its encoder initialized by the Fbk-to-DSU encoder. EncDec-Init is a DSU-Adapter model without the adapter layer.
+4.3 Training and Inference
+It is worth noting that we do not use extra audio data, e.g., Libri-Light Kahn et al. (2020) in our (pretraining) experiments. Furthermore, we apply the following conditions in (pre-)training:
+  • We skip training data that are longer than $30$ seconds (audio) or $1,024$ target tokens.
+  • We apply SpecAugment Park et al. (2019) with parameters: {$F=30,T=40,m_{F}=2,m_{T}=2$} on Filterbank inputs.
+  • We share the embedding weights when using a joint vocabulary in the DSU-to-Trl model.
+  • We set $\lambda_{\alpha}$ and $\lambda_{\beta}$ in CTC to $0.3$ and the smoothing parameter to $0.1$
+  • We initialize the encoder (decoder) with the last (best) checkpoint from the PT model.
+  • We use Adam optimizer with inverse square root scheduler for all model training.
+  • In all Fbk-to-Token models, the effective mini-batch size, warm-up steps, peak learning rate and training steps are $32$K frames, $25$K, $2$e33-3- 3 and $60$K steps respectively.
+  • Similarly, in all DSU-to-Trl models, we use $80$K tokens, $10$K, $5$e44-4- 4 and $50$K steps.
+  • Similarly, in Hu-Transformer, we use $4$M frames, $4$K, $1$e44-4- 4 and $300$K steps.
+In inference, we average the last 5 checkpoints and use beam size of $5$ in generation.
+All experiments are run on Nvidia A100 GPUs.
+It takes about $1$ day for $2$ A100 ($40$GB) GPUs to complete an experiment that uses Filterbank as speech inputs.
+AST model (#Params) BLEU chrF COMET-22-DA
+High Mid Low All High Mid Low All High Mid Low All
+Scratch (52M) $19.4$ $7.91$ $0.73$ $5.99$ $43.6$ $27.2$ $14.6$ $23.1$ $0.605$ $0.498$ $0.433$ $0.481$
+ASR-Pretraining (52M) $26.5$ $12.2$ $1.82$ $9.00$ $51.9$ $32.8$ $16.4$ $27.1$ $0.680$ $0.537$ $0.443$ $0.511$
+Hu-Transformer (113M) $24.3$ $11.4$ $2.18$ $8.60$ $49.9$ $31.9$ $17.0$ $26.8$ $0.650$ $0.522$ $0.439$ $0.499$
+DSU-Adapter (48M) $26.5$ $12.9$ $1.76$ $9.13$ $52.1$ $33.9$ $16.5$ $27.4$ $0.681$ $0.548$ $0.442$ $0.513$
+Table 1: Results in BLEU, chrF and COMET-22-DA on the test set of CoVoST-2 (X-En) by resource group.
+In all metrics, DSU-Adapter is much better than Hu-Transformer, which is $2.3$ times larger, in both "High" and "Mid" groups.
+DSU-Adapter, which does not requires transcripts in training, is also on a par with ASR-Pretraining.
+The best result in each group is denoted by ’_’.
+5 Results and Analysis
+Before discussing the results, it is worth noting that (1) Hu-Transformer is not memory-adaptive, and (2) ASR-Pretraining requires transcripts, unlike DSU which is self-supervised.
+Both methods are introduced for reference purposes of if such resources are available.
+5.1 Improvement brought by DSU-Adapter
+We divide the 21 language pairs by resource level into: 1) "High": {ca, de, es, fr}, 2) "Mid": {fa, it, pt, ru and zh}, 3) "Low": {ar, cy, et, id, ja, lv, mn, nl, sl, sv, ta, tr} and 4) "All": the 21 languages pairs.
+We report the average BLEU444nrefs:1—case:mixed—eff:no—tok:13a—smooth:exp—version:2.3.1 and chrF555nrefs:1—case:mixed—eff:yes—nc:6—nw:0—space:no—version:2.3.1 over the test sets of each group using sacrebleu Post (2018).
+In addition, we also provide the result in WMT22-COMET-DA Rei et al.
+(2022), which the source inputs are the gold-reference transcripts.
+Table 1 compares our DSU-Adapter and the baselines.
+Our DSU-Adapter is $3$ BLEU (in the group "All") higher than the Scratch model.
+This shows that our proposed method of using DSU-pretraining can strengthen direct end-to-end ST without requiring transcripts and remain flexible in memory footprint (smaller in size than the HuBERT model).
+Furthermore, it is better than Hu-Transformer in spite of having half the parameters.
+For "Mid" and "High", the improvement in BLEU is $1.49$ and $2.23$ points respectively, but it falls short by $0.42$ points for "Low".
+We also compare to ASR pretraining, which is not always applicable, e.g., in low-resource setting or perhaps even in an unwritten language Zhang et al.
+(2022a).
+Surprisingly, our adapter is on a par with it, and its BLEU is $0.13$ points better.
+The result remains consistent when it is measured in chrF and COMET.
+5.1.1 Language-specific performance
+Figure 3 shows the performance on each language pair in BLEU, chrF and COMET-22-DA.
+Our DSU-Adapter (in green triangles) show consistent improvement over the Scratch model (in blue circles) in all language pairs.
+Such improvement is rather surprising since HuBERT-Base was trained solely on English audio data.
+We hypothesized that the cross-lingual improvement is related to HuBERT’s ability to capture language independent features, e.g.
+phonetic properties Pasad et al.
+(2023).
+Compared with Hu-Transformer, DSU-Adapter maintains an evident improvement over most language pairs in both "High" and "Mid" groups.
+Exceptions are in "fa" and "pt", but the lags are almost negligible.
+In group "Low", Hu-Transformer is slightly better, especially in "nl" and "sv" pairs.
+However, most translation in this group is barely around $2$ BLEU, and the lags are small.
+In most language pairs, DSU-Adapter performs similarly to ASR-pretraining (in red diamonds), except translating from "ru" audios.
+The improvement in this "ru-en" pair makes DSU-Adapter to have an evident advantage of $0.7$ BLEU in the group "Mid".
 Has BPE $|\mathcal{V}|$ DSU Length DSU-to-Trl Enc-Init EncDec-Init DSU-Adapter
-on DSU? Length Ratio $20$M to $27$M $52$M to $70$M $46$M to $64$M $48$M to $67$M
+on DSU? Length Ratio ($20$M to $27$M) ($52$M to $70$M) ($46$M to $64$M) ($48$M to $67$M)
 No $1$K-$8$K $176$ $12.9$ $6.73$ $7.70$ $7.87$ $8.54$
@@
 $32$K $118$ $7.8$ $5.07$ $8.30$ $8.44$ $8.70$
-Table 2: Tokenization effect on 4 types of ST methods.
+Table 2: (DSU) tokenization effect on 4 ST methods.
+Each ST model’s performance on the CoVoST-2 test set is measured by BLEU on group "All".
+All 4 methods could perform better than the Scratch model of $5.99$ BLEU as shown on Table 1.
 In general, darker (brighter) cells refer to weaker (stronger) models.
 The best two models apply both BPE on the DSU and separate vocabularies in PT (cells in yellow).
-4 Experiments
-4.1 Data Preprocessing
-We follow standard practices to preprocess the CoVoST-2 X-En data, i.e., log-Mel Filterbank of $80$ dimensions for speech inputs and BPE with $8$K vocabulary on the English target, $16$K on the transcripts and $32$K on the DSU, unless otherwise specified.
-We use an English HuBERT-Base model to extract the DSU by first downsampling the CoVoST-2 audio to $16$KHz.
-Each audio data utterance is then converted into the DSU, i.e., the clustering indexes, by applying K-Means clustering (K=$1,000$) on its HuBERT representation from the 6th layer Lakhotia et al.
-(2021).
-We sample training data from each language pair to form a subset of $98$K instances to train the K-Means model.
-Further details are in Appendix A.1.
-4.2 Model Configuration
-All models are based on Transformer Vaswani et al.
-(2017) with implementations from fairseq Ott et al.
-(2019); Wang et al.
-(2020).
-In the Fbk-to-token [^1] Token refers to either transcriptions, DSU, or translations.
-models, the encoder has convolutional layers to downsample the Fbk by a factor of 4.
-There are 12-6 layers in the transformer encoder-decoder, whereas the embedding and feed-forward network (FFN) dimensions are 256 and 4,096 respectively, unless otherwise specified.
-It is worth noting that: (1) The Fbk-to-DSU model is not trained on the translations, so it is not directly comparable to the ST models.
-Its effect to ST lies on its pretrained encoder (Table 2).
-(2) The DSU-to-Trl model is a ST model, and it can also be used for initialization.
-Scratch is a ST model trained on the paired speech-translation data without pretraining.
-ASR Pretraining refers to a ST model whose encoder is initialized by a speech recognition task with CTC regularisation on the transcripts.
-DSU-to-Trl follows the Transformer used in text translation.
-We use 6-6 layers in the encoder-decoder which the dimension of embedding and FFN is 256 and 2,048 respectively.
-Despite its smaller model size, its inference requires storing and calling both HuBERT and K-Means models.
-Hu-Transformer uses the entire HuBERT as the speech encoder initialization Fang and Feng (2023). For comparison to our DSU-Adapter, its subsequent encoder-decoder also has 1-6 layers.
-DSU-Adapter is our proposed method.
-To better align the two pre-trained components, we also experiment with adding an extra encoder layer [^2] All parameters are trainable in ST finetuning.
-after the pre-trained encoder.
-Since its decoder is initialized by the DSU-to-Trl method, its decoder FFN dimension is $2,048$.
-Enc-Init is a ST model that has its encoder initialized by the Fbk-to-DSU encoder. EncDec-Init is a DSU-Adapter model without the adapter layer.
-4.3 Training and Inference
-We leave the details in Appendix A.3. It is worth noting that we do not use extra audio data, e.g., Libri-Light Kahn et al. (2020) in pretraining.
-5 Results and Analysis
-We divide the 21 language pairs by resource level into: 1) "High": {ca, de, es, fr}, 2) "Mid": {fa, it, pt, ru and zh}, 3) "Low": {ar, cy, et, id, ja, lv, mn, nl, sl, sv, ta, tr} and 4)"All": the 21 languages pairs.
-We report the average BLEU over the test sets of each group using sacrebleu Post (2018)333nrefs:1—case:mixed—eff:no—tok:13a—smooth:exp—version:2.3.1.
-Table 1 compares our DSU-Adapter and the baselines.
-Our DSU-Adapter is $3$ BLEU points (in the group "All") higher than Scratch.
-Furthermore, it is better than Hu-Transformer in spite of having half the parameters.
-For "Mid" and "High", the improvement is $1.49$ and $2.23$ points respectively, but it falls short by $0.42$ points for "Low".
-We also compare to ASR pretraining, which is not always applicable, e.g., in low-resource setting or perhaps even in an unwritten language Zhang et al.
-(2022a).
-Surprisingly, our adapter is on a par with it, i.e., $0.13$ points better.
-The result remains consistent when measured in chrF Popović (2015) and COMET [^4] WMT22-COMET-DA Rei et al.
-(2022), see Appendix A.4.
-5.1 Tokenization and its Transfer
-Table 2 presents the effect of tokenization.
-The 1st column indicates if BPE is applied on the DSU, and the 2nd column "$|\mathcal{V}|$" shows the vocabulary size, and if the model has a joint vocabulary or not in PT.
-For example, "$1$K-$8$K" means that we use a vocabulary of size $1$K for DSU and $8$K for English.
-These results are for all languages, please see Appendix A.4 for fine grained results.
-Effect of BPE on DSU-to-Trl
-When BPE is not applied on the DSU, the DSU-to-Trl models have a BLEU of $6.48$$\pm$ $0.26$ across the 6 vocabulary settings.
-Despite being smaller (<$30$M), these models are better than the Scratch model of $5.99$ BLEU.
-When BPE is applied, the DSU sequence length becomes a crucial factor.
-On one hand, it delivers the best model ("$8$K": $7.02$ BLEU).
-On the other hand, the result deteriorates a lot in the other 5 settings, i.e., a BLEU of $5.12$$\pm$ $0.83$.
-More importantly, BPE’s correlation with the model performance is not straightforward, neither is the length ratio.
-For an example, the "$32$K" model (DSU length of $118$) is about $2$ BLEU points behind to the "$8$K" model (DSU length of $150$).
-DSU-Adapter is more tokenization robust
-When BPE is applied on the DSU, DSU-Adapter has a higher BLEU of $8.86$$\pm$ $0.19$, in contrast to $8.17$$\pm$ $0.32$.
-This behaviour is opposed to the DSU-to-Trl model that has a lower BLEU of $5.54$$\pm$ $1.07$ (with also larger variance), in contrast to $6.48$$\pm$ $0.26$ when BPE is not used.
-The better scores with smaller variance show that DSU-Adapter is more tokenization robust.
-We see this as a benefit of introducing the DSU (SSL knowledge) via PT rather than as model inputs.
-Besides, using separate vocabularies in PT is preferred since it performs slightly better, and the DSU, which are not needed in the ST output, would not occupy the target vocabulary.
-Ablation
-Having similar size, DSU-Adapter is better than EncDec-Init and Enc-Init.
-The BLEU over their 12 settings is $8.51$$\pm$ $0.44$, $8.22$$\pm$ $0.51$, and $7.99$$\pm$ $0.61$ respectively.
-Encoder-initialization seems more crucial than decoder-initialization, as reflected by the fact that the best DSU-Adapter comes from a combination with the weakest DSU-to-Trl model of $4.43$ BLEU.
-5.2 CTC regularisation helps
+5.2 Tokenization effect to the DSU-to-Trl method and the DSU-Adapter method
+In this section, we investigate how tokenization, including BPE, affects the DSU-to-Trl method and the DSU-Adapter method.
+We are particularly interested in their robustness toward the tokenization, especially using BPE on the DSU, since tuning the quantization process and retraining the subsequent models is computationally expensive.
+In Table 2, the 1st column "Has BPE on DSU?" indicates if BPE is applied on the DSU.
+If "Yes", multiple DSU could be merged into one subword unit, e.g., ’#1 #456 #23 #999’ could be merged into ’#1#456#23#999’.
+The 2nd column "$|\mathcal{V}|$" shows the vocabulary configuration: its size, and if the model has a joint vocabulary.
+For example, "$1$K-$8$K" means that we use a vocabulary of size $1$K for DSU and a second vocabulary of size $8$K for English so that the DSU-to-Trl model would have separate vocabularies for the source (DSU) and target (English) sides.
+All results are in BLEU averaged over all language pairs, i.e., group "All".
+5.2.1 DSU-to-Trl: robust to tokenization?
+When BPE is not applied on the DSU, those 6 DSU-to-Trl models have $6.48$$\pm$ $0.26$ BLEU. Despite having smaller model size (<$30$M), they are better than the Scratch model of $5.99$ BLEU.
+When BPE is applied, the sequence length of DSU (DSU Length) could be shortened, which could in turn improve the performance, e.g.
+the best DSU-to-Trl model happens at configuration "$8$K" with $7.02$ BLEU.
+However, the DSU-to-Trl method is quite unstable to the use of BPE, as reflected by the $5.12$$\pm$ $0.83$ BLEU in the other 5 configurations.
+The correlation between the DSU sequence length, the source-target length ratio, and the ST performance is also not straightforward.
+For an example, the "$32$K" model (DSU length of $118$) is about $2$ BLEU behind to the "$8$K" model (DSU length of $150$).
+Therefore, applying BPE on the DSU for length reduction should remain cautious.
+5.2.2 The DSU-Adapter is more robust
+Unlike DSU-to-Trl method, DSU-Adapter benefits more when BPE is applied to the DSU.
+Our proposed method has $8.86$$\pm$ $0.19$ BLEU (over the 6 corresponding configurations), as opposed to $8.17$$\pm$ $0.32$ BLEU when BPE is not applied.
+This observation is opposed to the DSU-to-Trl method which only scores $5.54$$\pm$ $1.07$ BLEU (with also larger variance) when BPE is applied on the DSU but $6.48$$\pm$ $0.26$ when BPE is not used.
+The improved mean score and its smaller variance suggests that the DSU-Adapter method is more (DSU) tokenization robust.
+We see this as a benefit of introducing the DSU, i.e., the SSS model knowledge, via PT rather than as model inputs.
+On top of applying BPE on the DSU, using separate vocabularies in PT is preferred (the two yellow cells on Table 2) since it performs slightly better, and the DSU, which are not needed in the ST output, would not occupy the target vocabulary.
+5.2.3 Ablation: initialisation in DSU-Adapter
+Having similar model sizes, e.g.
+about $50$M parameters (Table 2), DSU-Adapter is better than both EncDec-Init and Enc-Init methods.
+The translation performance in BLEU (averaged over the 12 vocabularies) is $8.51$$\pm$ $0.44$, $8.22$$\pm$ $0.51$, and $7.99$$\pm$ $0.61$ respectively.
+Encoder-initialization seems more crucial than decoder-initialization, as reflected by the fact that the best DSU-Adapter model comes from a combination with the weakest DSU-to-Trl model of $4.43$ BLEU.
+5.3 Is CTC applicable also to DSU?
 Has CTC in High Mid Low All
@@
 Yes Yes $26.12$ $11.53$ $1.73$ $8.74$
-Table 3: Effect of CTC on Fbk-to-DSU PT and/or ST FT to the DSU-Adapter method.
-Table 3555Owing to the large number of configurations, we can experiment on 4 vocabularies only and then report their average.
-presents an analysis of using CTC in the Fbk-to-DSU PT and/or in the ST FT using the DSU-Adapter method.
-We find that CTC helps on either stage, but the gain is 0.27 BLEU more in finetuning.
-Using them jointly still helps, but the gain becomes marginal, i.e, 0.03 BLEU.
-6 Conclusion
-We propose to distill knowledge from the large SSL model via pretraining on its DSU to make ST models more compact for on-device deployment.
-Our compact model, i.e., the DSU-Adapter, shows stronger result over finetuning the corresponding (larger) HuBERT in multilingual ST task.
-In contrast to using DSU as model input, our method requires less components in inference and shows stronger robustness over the tokenization.
-Finally, our method is suitable for the low-resource setting since it requires no transcripts.
-7 Limitations
-In this work, we propose to distill knowledge from the informative SSL representations via DSU pretraining to make ST models more compact.
-In addition, we also discuss advantages of our method over using DSU as model inputs.
-In spite of its noticeable benefits, there are several factors that are not thoroughly explored:
+Table 3: Effect of CTC on Fbk-to-DSU PT and/or ST FT to the DSU-Adapter method. All results are in BLEU and the best in each group is denoted by ’_’.
+Similar to ST methods that use pretrained components, our method could be limited by the pretraining modality gap Liu et al.
+(2020); Le et al.
+(2023).
+Motivated by prior works, we investigate mitigating it with CTC.
+A crucial difference to the prior works is that our method uses DSU for pre-training rather than transcripts.
+We thus study applying CTC in our method at different training stages.
+Owing to the large number of vocabulary configurations on Table 2, we only experiment with: 1) "No-BPE $1$K-$8$K", 2) "BPE $8$K", 3) "BPE $32$K" and 4) "BPE $32$K-$8$K".
+In each training stage, we report the effect of CTC to the ST performance (per resource group) by averaging the BLEU of these $4$ configurations.
+Table 3 presents the analysis of applying CTC on our DSU-Adapter method.
+The training condition "Has CTC in DSU PT" refers to the case of applying CTC on the discrete speech units in Fbk-to-DSU pretraining, whereas "Has CTC in ST FT" refers to the case of applying CTC on the translations in ST finetuning, i.e., on the paired Fbk-Trl data.
+Our result shows that CTC helps on either stage, but the gain is $0.27$ BLEU more in ST finetuning.
+Using them jointly still helps, but the marginal gain is barely $0.03$ BLEU.
+6 Limitations and future works
+In the previous sections, we discuss the noticeable benefits of our DSU-pretraining method in creating a more compact ST model.
+In spite of this, there are several factors that are not thoroughly explored and could improve the model performance further:
 K-Means clustering
 We did not inspect the clustering size (fixed to $1,000$) and the number of training instances (only fixed to $98,000$) used in training the K-Means clustering model.
+Apart from tuning its hyper-parameters, using other techniques, such as residual vector quantisation Zeghidour et al.
+(2021); Défossez et al.
+(2022) and multiple codebooks Guo et al.
+(2023), might bring better improvement.
 Other acoustic encoders
@@
 This stronger encoders should provide further gains for our method since they also enjoy the benefit of pretraining.
-Self-Supervised Models
-We did not experiment using other layers (only fixed to the 6th layer) from the HuBERT model nor from stronger SSL models to extract the DSU.
-These two factors potentially increase the DSU quality and potentially provide better improvement for our method while maintaining the same model size.
+A stronger pretrained decoder
+Apart from strengthening the encoder, the DSU-to-Trl model and hence its decoder (used in initialisation) could also be improved, e.g.
+via back-translation, up-sampling the textual sequence Yan et al.
+(2024) and pretraining with more text data, while maintaining the small decoder size.
+Further analyses
+In addition to improving our pretraining method for better model compactness, there are other related research directions worth further analyzing.
+One direction would be how, in terms of acoustic pretraining, DSU compared with transcripts (if available in that language) over different data scales.
+Another interesting research direction would be the comparison and analysis of using DSU or dense features in a large pretrained model setting, such as Whisper Radford et al.
+(2023) and Large Language Models.
+7 Conclusion
+In this paper, we consider a memory-constrained setting for ST.
+Our proposed method uses DSU in the form of pretraining to distil the knowledge from the Self-Supervised Speech model to creating more compact Speech-to-text Translation.
+Our compact model, i.e., the DSU-Adapter, shows strong and consistent improvements in three evaluation metrics over the baselines.
+In contrast to using DSU as model inputs, our method does not require quantization modules in inference and shows stronger robustness to the DSU tokenization.
+Finally, our method requires no transcripts, making it also suitable for low-resource setting.
+Figure 3: Results in BLEU, chrF and COMET-22-DA on each language pair of CoVoST-2 (X-En).
 Acknowledgements
@@
 Baskerville was funded by the EPSRC and UKRI through the World Class Labs scheme (EP/T022221/1) and the Digital Research Infrastructure programme (EP/W032244/1) and is operated by Advanced Research Computing at the University of Birmingham.
-Appendix A Appendix
-A.1 Data preprocessing
-log Mel Filterbank
-The 80-D log Mel Filterbank is computed every 10ms with a 25ms window, followed by feature normalisation using statistics, i.e., mean and variance, computed over each channel.
-BPE
-We use the BPE implementation from sentencepiece Kudo and Richardson (2018) and obtain vocabulary of size 8K on the English target, 16K on the transcripts and 32K on the DSU, unless otherwise specified.
-HuBERT
-K-Means Clustering
-We use MiniBatchKMeans from sklearn with a clustering size of $1,000$ and default settings in other hyper-parameters.
-When training the clustering model, we divide the 21 language pairs into three groups: 1) {ar, cy, et, id, ja, lv, mn, sl, sv, ta, tr}, 2) {nl, pt, ru, zh} and 3) {ca, de, es, fa, fr, it} for sampling.
-For each language pair in group 1), we sample $1$K instances, whereas it becomes $3$K in group 2) and $12.5$K in group 3), resulting in a total of $98$K instances.
-A.2 Model Configuration
-We use "pre" layer-normalization Nguyen and Salazar (2019) in DSU-to-Trl.
-A.3 Training and Inference
-In training,
-  • we skip training data that are more than $3$K frames for Filterbank, $320$K frames for HuBERT input or $1,024$ tokens.
-  • we apply SpecAugment Park et al. (2019) with parameters: {$F=30,T=40,m_{F}=2,m_{T}=2$} on Filterbank inputs.
-  • we share the embedding weights when using a joint vocabulary in the DSU-to-Trl model.
-  • we set $\lambda_{\alpha}$ and $\lambda_{\beta}$ in CTC to $0.3$ and the smoothing parameter in LS to $0.1$
-  • we initialize the encoder (decoder) with the last (best) checkpoint from the PT model.
-All models are optimized by Adam Kingma and Ba (2015) with inverse square root scheduler.
-For simplicity, we provide major training parameters for each model in the format of {effective batch size, warm-up steps, peak learning rate, maximum training steps}:
-Fbk-to-Token:
-{32K frames, 25K, 2e-3, 60K}
-DSU-to-Trl:
-{80K tokens, 10K, 5e-4, 50K}
-Hu-Transformer:
-{4M frames, 4K, 1e-4, 300K}
-In inference, we average the last 5 checkpoints and use beam size of 5 in generation.
-All experiments are run on Nvidia A100 GPUs.
-It takes about 1 day for 2 A100 (40GB) GPUs to complete an experiment that uses Filterbank as inputs.
-A.4 Results and Analysis
-In addition to BLEU (see Table 8 for the result of each language), we also provide the result of Table 1 using chrF777nrefs:1—case:mixed—eff:yes—nc:6—nw:0—space:no—version:2.3.1 and COMET.
-In order to compute the COMET score, we feed the gold-reference transcripts as the source sentence.
-Table 4 and Table 6 show the average result in chrF and COMET respectively over the four groups defined in section 5.
-Their results follow the findings in BLEU: the DSU-Adapter is better than Hu-Transformer and on a par with ASR pretraining.
-Further details, such as training data size (in terms of the number of utterances rather than hours888”zh” has 10 hours of audio data, whereas ”nl” has only 7 hours.
-Therefore, ”zh” is classified into ”Mid” resource group even though it has slightly less number of utterances than ”nl”.) and per language performance, are shown on Table 5 for chrF and Table 7 for COMET.
-CTC Regularisation
-On Table 3, we average the results over this four configurations: 1) "No-BPE 1K-8K", 2) "BPE 8K", 3) "BPE 32K" and 4) "BPE 32K-8K".
-AST model Transcripts Model size High Mid Low All
-Scratch No 52M $43.58$ $27.22$ $14.62$ $23.10$
-ASR Pretraining Yes 52M $51.90$ $32.82$ $16.38$ $27.10$
-Hu-Transformer No 113M $49.88$ $31.94$ $16.96$ $26.80$
-DSU-Adapter No 48M $52.08$ $33.94$ $16.46$ $27.40$
-Table 4: Results of Table 1 in chrF score by resource group.
-Similar to the findings in BLEU, DSU-Adapter is far better than Hu-Transformer, in particular for "High" and "Mid", and on a par with ASR Pretraining.
-The best result in each group is denoted by ’_’.
-Languages #Training data Resources Scratch ASR-PT Hu-Transformer DSU-Adapter
-ca 95,854 High $42.5$ $50.9$ $49.8$ $51.5$
-de 127,824 $35.4$ $45.1$ $43.6$ $45.4$
-es 79,013 $48.1$ $56$ $52.8$ $55.7$
-fr 207,372 $48.3$ $55.6$ $53.3$ $55.7$
-fa 53,949 Mid $15.6$ $16.9$ $19.9$ $16.9$
-it 31,698 $40.6$ $50.2$ $46.4$ $50.3$
-pt 9,158 $25$ $31.3$ $33.8$ $32.3$
-ru 12,112 $35.6$ $42.6$ $40$ $47.7$
-zh 7,085 $19.3$ $23.1$ $19.6$ $22.5$
-ar 2,283 Low $12.5$ $14.4$ $13.8$ $13.2$
-cy 1,241 $14.5$ $16.3$ $15.8$ $15.4$
-et 1,782 $17.9$ $21.3$ $19$ $19.9$
-id 1,243 $13.8$ $14.2$ $14.9$ $14.5$
-ja 1,119 $14$ $14$ $13.7$ $14.4$
-lv 2,337 $13.9$ $14.5$ $15.9$ $15.4$
-mn 2,067 $14.5$ $15.4$ $14.3$ $14.8$
-nl 7,108 $17.8$ $24.3$ $30.8$ $27$
-sl 1,843 $13.9$ $15.4$ $15.3$ $15.4$
-sv 2,160 $13.2$ $13.8$ $17.3$ $14.5$
-ta 1,358 $12$ $11.9$ $12$ $12.3$
-tr 3,966 $17.4$ $21.1$ $20.7$ $20.7$
-Table 5: Results of Table 1 in chrF score for each language pair. The best result in each group is denoted by ’_’.
-AST model Transcripts Model size High Mid Low All
-Scratch No 52M $0.605$ $0.498$ $0.433$ $0.481$
-ASR Pretraining Yes 52M $0.680$ $0.537$ $0.443$ $0.511$
-Hu-Transformer No 113M $0.650$ $0.522$ $0.439$ $0.499$
-DSU-Adapter No 48M $0.681$ $0.548$ $0.442$ $0.513$
-Table 6: Results of Table 1 in COMET score by resource group.
-Similar to the findings in BLEU, DSU-Adapter is far better than Hu-Transformer, in particular for "High" and "Mid", and on a par with ASR Pretraining.
-The best result in each group is denoted by ’_’.
-Languages #Training data Resources Scratch ASR-PT Hu-Transformer DSU-Adapter
-ca 95,854 High $0.607$ $0.681$ $0.654$ $0.684$
-de 127,824 $0.535$ $0.608$ $0.588$ $0.613$
-es 79,013 $0.620$ $0.696$ $0.654$ $0.692$
-fr 207,372 $0.660$ $0.736$ $0.704$ $0.735$
-fa 53,949 Mid $0.479$ $0.494$ $0.501$ $0.493$
-it 31,698 $0.557$ $0.640$ $0.594$ $0.641$
-pt 9,158 $0.468$ $0.497$ $0.510$ $0.508$
-ru 12,112 $0.551$ $0.580$ $0.563$ $0.630$
-zh 7,085 $0.434$ $0.472$ $0.440$ $0.467$
-ar 2,283 Low $0.438$ $0.465$ $0.45$ $0.445$
-cy 1,241 $0.420$ $0.448$ $0.424$ $0.426$
-et 1,782 $0.413$ $0.432$ $0.407$ $0.423$
-id 1,243 $0.420$ $0.406$ $0.416$ $0.412$
-ja 1,119 $0.417$ $0.412$ $0.411$ $0.422$
-lv 2,337 $0.457$ $0.453$ $0.465$ $0.461$
-mn 2,067 $0.429$ $0.441$ $0.419$ $0.437$
-nl 7,108 $0.427$ $0.460$ $0.483$ 0.477
-sl 1,843 $0.430$ $0.428$ $0.433$ $0.428$
-sv 2,160 $0.454$ $0.448$ $0.472$ $0.454$
-ta 1,358 $0.413$ $0.406$ $0.403$ $0.411$
-tr 3,966 $0.480$ $0.519$ $0.484$ $0.514$
-Table 7: Results of Table 1 in COMET score for each language pair. The best result in each group is denoted by ’_’.
-Languages #Training data Resources Scratch ASR-PT Hu-Transformer DSU-Adapter
-ca 95,854 High $18.24$ $25.66$ $24.24$ $25.98$
-de 127,824 $13.97$ $21.06$ $19.28$ $21.32$
-es 79,013 $22.37$ $29.8$ $26.64$ $29.34$
-fr 207,372 $22.91$ $29.6$ $27.09$ $29.52$
-fa 53,949 Mid $1.9$ $3.07$ $4.63$ $2.84$
-it 31,698 $15.1$ $23.55$ $19.98$ $23.36$
-pt 9,158 $7.01$ $11.2$ $13.13$ $11.69$
-ru 12,112 $11.97$ $16.87$ $14.95$ $21$
-zh 7,085 $3.58$ $6.39$ $4.4$ $5.67$
-ar 2,283 Low $0.44$ $1.96$ $1.26$ $0.48$
-cy 1,241 $0.64$ $1.24$ $1.21$ $0.72$
-et 1,782 $0.49$ $1.46$ $0.61$ $0.93$
-id 1,243 $0.31$ $0.67$ $1.27$ $0.79$
-ja 1,119 $0.27$ $0.67$ $0.36$ $0.62$
-lv 2,337 $1.22$ $1.65$ $2.63$ $2.06$
-mn 2,067 $0.16$ $0.44$ $0.31$ $0.19$
-nl 7,108 $2.14$ $6.95$ $10.75$ $8.46$
-sl 1,843 $0.41$ $1.83$ $1.23$ $1.81$
-sv 2,160 $0.9$ $1.28$ $3.4$ $1.76$
-ta 1,358 $0.14$ $0.14$ $0.13$ $0.12$
-tr 3,966 $1.67$ $3.5$ $3.04$ $3.14$
-Table 8: Results of Table 1 in BLEU score for each language pair. The best result in each group is denoted by ’_’.